### v2

@@ 2 A relative-type inexact pALM
In this section, we focus on developing a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). The algorithmic framework is developed based on the parametric convex duality framework (see, for example, [39, 40] and [43, Chapter 11]).
@@
Note that $G$ is proper closed convex since $f^{*}$ is proper closed convex. Then, the (ordinary) Lagrangian function of problem (1.2) can be defined by taking the concave conjugate of $G$ with respect to its second argument (see [43, Definition 11.45]), that is,
@@
Clearly, $\ell$ is convex in its first argument and concave in the second argument. For a given penalty parameter $\sigma>0$, the augmented Lagrangian function of problem (1.2) is defined by (see [43, Example 11.57])
@@
Compared to the recent semismooth Newton based inexact proximal augmented Lagrangian (Snipal) method developed in [24, Section 3], our ripALM in Algorithm 1 employs a significantly different error criterion (2.5) for solving the subproblem (2.4). Specifically, in our context, Snipal requires the error term $\Delta^{k+1}$ to satisfy
@@
with $\Delta_{x}^{k+1}:=\bm{x}^{k+1}-\bm{x}^{k}$ and $\Delta_{y}^{k+1}:=\bm{y}^{k+1}-\bm{y}^{k}$, to guarantee an asymptotic (super)linear convergence rate. [^3] Note that the error criterion (A) in (2.7) alone is sufficient for establishing the global convergence of the Snipal; see [24, Section 3]. Both error criteria (A) and (B) in (2.7) are of the absolute type, meaning that they require the pre-specification of two summable tolerance parameter sequences, $\{\varepsilon_{k}\}\subseteq[0,\infty)$ and $\{\delta_{k}\}\subseteq[0,1)$, to control the error incurred in the inexact minimization of the subproblem. Since there is generally no direct guidance on optimally selecting these tolerance parameters to achieve good convergence efficiency, this absolute-type criterion typically requires careful tuning for the tolerance parameters, which may make Snipal less user-friendly in practical implementations. In contrast, the error criterion (2.5) used in ripALM is of the relative type, meaning that the error $2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\big\|\sigma_{k}\Delta^{k+1}\big\|^{2}$ is regulated by a tentative successive difference related to the progress of the algorithm. Moreover, this relative-type criterion only involves a single tolerance parameter $\rho\in[0,1)$, thus simplifying the process of parameter tuning in both computation and implementation, as we shall see in subsection 5.1.
Thanks to the advantage of eliminating the need to select an infinite sequence of tolerance parameters, relative-type error criteria have been widely adopted in numerous well-known algorithms (e.g., PPA, ALM, ADMM, FISTA, etc.) to approximately solve subproblems over the past two decades. This trend began with the seminal works of Solodov and Svaiter [45, 46, 47, 48, 49], which has since inspired the development of numerous relative-type inexact algorithms; see, for example, [4, 14, 15, 33, 56, 58, 59]. Following this line of research, Yang et al. [58] recently developed a corrected inexact proximal augmented Lagrangian method (ciPALM) for solving a class of group-quadratic regularized optimal transport problems. The error criterion used there can be described as follows: choose $\rho\in[0,1)$, at each iteration, approximately solve the subproblem (2.4) to find a point $\widetilde{\bm{y}}^{k+1}$ such that
@@
When (2.8) is satisfied, the multiplier is updated as usual: $\bm{x}^{k+1}=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big)$. This is then followed by an extra correction step: $\bm{y}^{k+1}=\bm{y}^{k}-\tau_{k}^{-1}\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right)$. After these updates, the algorithm proceeds to the next iteration. One can see that the error criterion (2.8) used in ciPALM is also of the relative type and looks even simpler than (2.5) used in our ripALM. However, we should point out that ciPALM requires an additional correction step to ensure its convergence, and hence it may only be viewed as a pALM-like algorithm. Moreover, our numerical results indicate that this correction step tends to degrade the practical performance of the pALM framework. This raises the natural question of whether one can design a relative-type error criterion directly for the vanilla pALM framework. Addressing this question serves as the main motivation of this work. To this end, we propose the relative error criterion (2.5), where the error variable $\bm{w}^{k}$ is used solely in the construction of the error criterion (2.5) and does not directly influence either the objective function in (2.4) or the updating rules of the primal and dual variables. Our numerical results show that the proposed ripALM may offer greater robustness and efficiency; see subsection 5.1 for numerical comparisons.
@@
Our relative error criterion (2.5) is inspired by Eckstein and Silva’s practical relative error criterion [14], which was developed for the approximate minimization of the subproblems in the vanilla ALM (i.e., without the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4) in our context). Unlike their work, we suggest incorporating $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4). This proximal term guarantees the existence and uniqueness of the optimal solution of the strongly convex subproblem (2.4). Introducing the proximal term also ensures the positive definiteness of the generalized Hessian/Jacobian of the objective function in (2.4), thereby facilitating the application of the semi-smooth Newton method to the subproblem (2.4) effectively, as shown in Section 4.1. In addition, as we will see later in the next section, our ripALM not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as Eckstein and Silva’s relative-type inexact ALM (denoted by ES-iALM) developed in [1, 14, 61], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker error bound assumption (than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin), the sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates.
@@ 3 Convergence analysis
Moreover, all the subgradient maps $\partial G$, $\partial F$, and $\partial\ell$ are maximal monotone operators, and satisfy that (see also [14, equation (23)]):
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bm{0}\in\partial f(\bm{x})-A^{\top}\bm{y},\quad A\bm{x}-\bm{b}=\bm{0},$
@@
Next, we study the asymptotic (super)linear convergence rate of our ripALM under an error bound condition presented in Assumption A. As noted in [24, Lemma 2.4], this error-bound condition is weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [61] to establish the asymptotic (super)linear convergence rate for Eckstein and Silva’s relative-type inexact ALM, while the former, weaker condition has also been employed in [24] and [58] to establish the asymptotic (super)linear convergence rate for Snipal and ciPALM, respectively.
@@
One can see from the expression of $\gamma_{k}$ in Theorem 3.2 that, after a finite number of iterations, $\gamma_{k}$ becomes proportional to the squared penalty parameter $\sigma_{k}^{2}$, provided that $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy the conditions in (3.5). Therefore, if $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$, the convergence factor $\mu_{k}$ can be less than 1 when $\sigma_{k}$ is sufficiently large, and it approaches 0 as $\sigma_{k}$ tends to $\infty$. This readily demonstrates that the sequence $\left\{(\bm{y}^{k},\bm{x}^{k})\right\}$ converges to the set of saddle points at a Q-(super)linear rate if $\sigma_{k}$ is sufficiently large and $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$. In practical implementations, one could simply choose an increasing sequence of $\{\sigma_{k}\}$ with $\sigma_{k}\uparrow+\infty$, and set $\tau_{\min}\geq 4$ to ensure that $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$ for any $\rho\in[0,1)$. In contrast, as discussed in [58, Remark 2.1], ciPALM (which is also of relative-type but uses the error criterion (2.8)) should require $\rho<\frac{1}{3}$ to guarantee an asymptotic (super)linear convergence rate under the same error bound condition in Assumption A. This highlights another advantage of our ripALM, as it offers greater flexibility in choosing the tolerance parameter $\rho$.
Before closing this section, we summarize in Table 1 a brief comparison among the related inexact (p)ALMs, based on the discussions in Section 2 and the convergence results established in this section. From the comparison, we see that ripALM enjoys stronger theoretical guarantees than ES-iALM and, unlike ciPALM, does not require correction steps.
Table 1: A brief comparison among ES-iALM, Snipal, ciPALM, and ripALM. The convergence results for each algorithm are presented in the context of solving problem (1.2), which is consistent with our setting. In the table, “Seq. Conv.”: sequential convergence; “Abs.”: absolute-type; “Rel.”: relative-type; ✔E: be convergent under error bound condition.
Methods Type
Vanilla
(p)ALM
Proximal
Term
Boundedness Seq. Conv.
Asymptotic
(super)linear Rate
(under Error Bound)
$\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$ $\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$
ES-iALM [1, 14, 61] Rel. $\{\bm{x}^{k}\}$
Snipal [24] Abs. $\{(\bm{y}^{k},\bm{x}^{k})\}$
ciPALM [58] Rel.
$\{(\bm{y}^{k},\bm{x}^{k})\}$
ripALM [this work] Rel. E $\{(\bm{y}^{k},\bm{x}^{k})\}$
@@ 4 Applications of ripALM
In this section, we present the application of the ripALM in Algorithm 1 to two important problems: the quadratically regularized optimal transport problem and the basis pursuit denoising problem.
@@ 4.1 Quadratically regularized optimal transport
Then, using similar arguments as in the proof of [58, Proposition 4.1], one can show that $\nabla\Psi(\cdot)$ is strongly semi-smooth with respect to $\widehat{\partial}(\nabla\Psi)(\cdot)$, and thus the Ssn method is applicable. More importantly, for any $(\bm{u},\bm{v})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, all elements of $\widehat{\partial}(\nabla\Psi)(\bm{u},\bm{v})$ are positive definite. This ensures the direct applicability of the Ssn method without the need for a specific regularity condition, such as the primal constraint nondegeneracy condition [62], highlighting the advantage of incorporating a proximal term. The detailed description of the Ssn method for solving equation (4.7) is presented in Algorithm 2. We refer readers to [22, Theorem 3.6] for its detailed convergence results.
@@ 4.2 Basis pursuit denoising
Using the relation $\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}(\bm{x})=\bm{x}-\mathtt{prox}_{\widehat{\kappa}\|\cdot\|_{2}}(\bm{x})$, Lemma 2.1 in [60], and similar arguments to those in [58, Proposition 4.1], one can show that $\nabla\Phi$ is strongly semismooth with respect to the multifunction $\widehat{\partial}(\nabla\Phi)$. This property enables the application of the Ssn method to solve (4.16). Since the algorithmic framework is identical to Algorithm 2, we omit the detailed description for simplicity. Moreover, note from (4.17) that, for any $\bm{y}\in\mathbb{R}^{m}$, all matrices in $\widehat{\partial}(\nabla\Phi)(\bm{y})$ are symmetric positive definite. This ensures the local superlinear (even quadratic) convergence rate of the Ssn method, without requiring additional regularity conditions. These favorable theoretical properties form the basis for the encouraging numerical performance, as observed from our numerical results in the next section.
@@
to obtain the Newton direction $\bm{d}$, where $H\in\widehat{\partial}(\nabla\Phi)(\bm{y})$ with $\widehat{\partial}(\nabla\Phi)$ defined in (4.17). Inspired by the techniques developed in [22, Section 3.3], we can exploit the special structure of $H$ to enhance the computational efficiency of solving (4.18). However, the presence of the additional block $V$ in $H$ introduces nontrivial implementation difficulties. Detailed discussions on efficiently solving (4.18) are provided in Appendix D.

### v3

@@ 2 A relative-type inexact pALM
In this section, we focus on developing a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). The algorithmic framework is developed based on the parametric convex duality framework (see, for example, [35, 36] and [34, Chapter 11]).
@@
Note that $G$ is proper closed convex since $f^{*}$ is proper closed convex. Then, the (ordinary) Lagrangian function of problem (1.2) can be defined by taking the concave conjugate of $G$ with respect to its second argument (see [34, Definition 11.45]), that is,
@@
Clearly, $\ell$ is convex in its first argument and concave in the second argument. For a given penalty parameter $\sigma>0$, the augmented Lagrangian function of problem (1.2) is defined by (see [34, Example 11.57])
@@
Compared to the recent semismooth Newton based inexact proximal augmented Lagrangian (Snipal) method developed in [19, Section 3], our ripALM in Algorithm 1 employs a significantly different error criterion (2.5) for solving the subproblem (2.4). Specifically, in our context, Snipal requires the error term $\Delta^{k+1}$ to satisfy
@@
with $\Delta_{x}^{k+1}:=\bm{x}^{k+1}-\bm{x}^{k}$ and $\Delta_{y}^{k+1}:=\bm{y}^{k+1}-\bm{y}^{k}$, to guarantee an asymptotic (super)linear convergence rate. [^2] Note that the error criterion (A) in (2.7) alone is sufficient for establishing the global convergence of the Snipal; see [19, Section 3]. Both error criteria (A) and (B) in (2.7) are of the absolute type, meaning that they require the pre-specification of two summable tolerance parameter sequences, $\{\varepsilon_{k}\}\subseteq[0,\infty)$ and $\{\delta_{k}\}\subseteq[0,1)$, to control the error incurred in the inexact minimization of the subproblem. Since there is generally no direct guidance on optimally selecting these tolerance parameters to achieve good convergence efficiency, this absolute-type criterion typically requires careful tuning for the tolerance parameters, which may make Snipal less user-friendly in practical implementations. In contrast, the error criterion (2.5) used in ripALM is of the relative type, meaning that the error $2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\big\|\sigma_{k}\Delta^{k+1}\big\|^{2}$ is regulated by a tentative successive difference related to the progress of the algorithm. Moreover, this relative-type criterion only involves a single tolerance parameter $\rho\in[0,1)$, thus simplifying the process of parameter tuning in both computation and implementation, as we shall see in Section 5.
Thanks to the advantage of eliminating the need to select an infinite sequence of tolerance parameters, relative-type error criteria have been widely adopted in numerous well-known algorithms (e.g., PPA, ALM, ADMM, FISTA, etc.) to approximately solve subproblems over the past two decades. This trend began with the seminal works of Solodov and Svaiter [40, 41, 42], which has since inspired the development of numerous relative-type inexact algorithms; see, for example, [4, 9, 10, 28, 46, 48, 49]. Following this line of research, Yang et al. [48] recently developed a corrected inexact proximal augmented Lagrangian method (ciPALM) for solving a class of group-quadratic regularized optimal transport problems. The error criterion used there can be described as follows: choose $\rho\in[0,1)$, at each iteration, approximately solve the subproblem (2.4) to find a point $\widetilde{\bm{y}}^{k+1}$ such that
@@
When (2.8) is satisfied, the multiplier is updated as usual: $\bm{x}^{k+1}=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big)$. This is then followed by an extra correction step: $\bm{y}^{k+1}=\bm{y}^{k}-\tau_{k}^{-1}\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right)$. After these updates, the algorithm proceeds to the next iteration. One can see that the error criterion (2.8) used in ciPALM is also of the relative type and looks even simpler than (2.5) used in our ripALM. However, we should point out that ciPALM requires an additional correction step to ensure its convergence, and hence it may only be viewed as a pALM-like algorithm. Moreover, our numerical results indicate that this correction step tends to degrade the practical performance of the pALM framework. This raises the natural question of whether one can design a relative-type error criterion directly for the vanilla pALM framework. Addressing this question serves as the main motivation of this work. To this end, we propose the relative error criterion (2.5), where the error variable $\bm{w}^{k}$ is used solely in the construction of the error criterion (2.5) and does not directly influence either the objective function in (2.4) or the updating rules of the primal and dual variables. Our numerical results show that the proposed ripALM may offer greater robustness and efficiency; see Section 5 for numerical comparisons.
@@
Our relative error criterion (2.5) is inspired by Eckstein and Silva’s practical relative error criterion [9], which was developed for the approximate minimization of the subproblems in the vanilla ALM (i.e., without the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4) in our context). Unlike their work, we suggest incorporating $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4). This proximal term guarantees the existence and uniqueness of the optimal solution of the strongly convex subproblem (2.4). Introducing the proximal term also ensures the positive definiteness of the generalized Hessian/Jacobian of the objective function in (2.4), thereby facilitating the effective application of the semi-smooth Newton method to the subproblem (2.4), as shown in Section 4.1. In addition, as we will see later in the next section, our ripALM not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as Eckstein and Silva’s relative-type inexact ALM (denoted by ES-iALM) developed in [1, 9, 51], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker error bound assumption (than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin), the sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates.
For ease of comparison, Table 1 summarizes the main algorithmic and theoretical differences between ripALM and the related methods discussed above. To the best of our knowledge, ripALM is the first relative-type inexact version of the vanilla pALM (1.4) that avoids both summable tolerance sequences and correction steps.
Table 1: A comparison of ES-iALM, Snipal, ciPALM, and ripALM for solving problem (1.2). In the table, “Seq. Conv.” means sequential convergence; “Abs.” means absolute-type; “Rel.” means relative-type; “✔E” indicates convergence under an error bound condition.
Methods Type Vanilla (p)ALM Proximal Term Boundedness Seq. Conv. Asymptotic (super)linear Rate (under Error Bound)
$\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$ $\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$
ES-iALM [1, 9, 51] Rel. $\{\bm{x}^{k}\}$
Snipal [19] Abs. $\{(\bm{y}^{k},\bm{x}^{k})\}$
ciPALM [48] Rel. $\{(\bm{y}^{k},\bm{x}^{k})\}$
ripALM [this work] Rel. E $\{(\bm{y}^{k},\bm{x}^{k})\}$
@@ 3 Convergence analysis
Moreover, all the subgradient maps $\partial G$, $\partial F$, and $\partial\ell$ are maximal monotone operators, and satisfy that (see also [9, equation (23)]):
@@
$\bm{0}\in\partial f(\bm{x})-A^{\top}\bm{y},\quad A\bm{x}-\bm{b}=\bm{0},$
@@
Next, we study the asymptotic (super)linear convergence rate of our ripALM under an error bound condition presented in Assumption A. As noted in [19, Lemma 2.4], this error bound condition is weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [51] to establish the asymptotic (super)linear convergence rate for Eckstein and Silva’s relative-type inexact ALM, while the former, weaker condition has also been employed in [19] and [48] to establish the asymptotic (super)linear convergence rate for Snipal and ciPALM, respectively.
@@
One can see from the expression of $\gamma_{k}$ in Theorem 3.2 that, after a finite number of iterations, $\gamma_{k}$ becomes proportional to the squared penalty parameter $\sigma_{k}^{2}$, provided that $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy the conditions in (3.5). Therefore, if $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$, the convergence factor $\mu_{k}$ can be less than 1 when $\sigma_{k}$ is sufficiently large, and it approaches 0 as $\sigma_{k}$ tends to $\infty$. This readily demonstrates that the sequence $\left\{(\bm{y}^{k},\bm{x}^{k})\right\}$ converges to the set of saddle points at a Q-(super)linear rate if $\sigma_{k}$ is sufficiently large and $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$. In practical implementations, one could simply choose an increasing sequence of $\{\sigma_{k}\}$ with $\sigma_{k}\uparrow+\infty$, and set $\tau_{\min}\geq 4$ to ensure that $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$ for any $\rho\in[0,1)$. In contrast, as discussed in [48, Remark 2.1], ciPALM (which is also of relative-type but uses the error criterion (2.8)) should require $\rho<\frac{1}{3}$ to guarantee an asymptotic (super)linear convergence rate under the same error bound condition in Assumption A. This highlights another advantage of our ripALM, as it offers greater flexibility in choosing the tolerance parameter $\rho$.
In summary, based on the discussions in Section 2 and the convergence results established in this section, we see that ripALM enjoys stronger theoretical guarantees than ES-iALM and, unlike ciPALM, does not require correction steps.
@@ 4 Applications of ripALM
In this section, we present applications of the ripALM in Algorithm 1 to two important problems: the quadratically regularized optimal transport problem and the basis pursuit denoising problem.
@@ 4.1 Quadratically regularized optimal transport
Then, using similar arguments as in the proof of [48, Proposition 4.1], one can show that $\nabla\Psi(\cdot)$ is strongly semi-smooth with respect to $\widehat{\partial}(\nabla\Psi)(\cdot)$, and thus the Ssn method is applicable. More importantly, for any $(\bm{u},\bm{v})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, all elements of $\widehat{\partial}(\nabla\Psi)(\bm{u},\bm{v})$ are positive definite. This ensures the direct applicability of the Ssn method without the need for a specific regularity condition, such as the primal constraint nondegeneracy condition [52], highlighting the advantage of incorporating a proximal term. The detailed description of the Ssn method for solving equation (4.7) is presented in Algorithm 2. We refer readers to [17, Theorem 3.6] for its detailed convergence results.
@@ 4.2 Basis pursuit denoising
Using the relation $\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}(\bm{x})=\bm{x}-\mathtt{prox}_{\widehat{\kappa}\|\cdot\|_{2}}(\bm{x})$, Lemma 2.1 in [50], and similar arguments to those in [48, Proposition 4.1], one can show that $\nabla\Phi$ is strongly semismooth with respect to the multifunction $\widehat{\partial}(\nabla\Phi)$. This property enables the application of the Ssn method to solve (4.16). Since the algorithmic framework is identical to Algorithm 2, we omit the detailed description for simplicity. Moreover, note from (4.17) that, for any $\bm{y}\in\mathbb{R}^{m}$, all matrices in $\widehat{\partial}(\nabla\Phi)(\bm{y})$ are symmetric positive definite. This ensures the local superlinear (even quadratic) convergence rate of the Ssn method, without requiring additional regularity conditions. These favorable theoretical properties form the basis for the encouraging numerical performance, as observed from our numerical results in the next section.
@@
to obtain the Newton direction $\bm{d}$, where $H\in\widehat{\partial}(\nabla\Phi)(\bm{y})$ with $\widehat{\partial}(\nabla\Phi)$ defined in (4.17). Inspired by the techniques developed in [17, Section 3.3], we can exploit the special structure of $H$ to enhance the computational efficiency of solving (4.18). However, the presence of the additional block $V$ in $H$ introduces nontrivial implementation difficulties. Detailed discussions on efficiently solving (4.18) are provided in Appendix D.